Matt and Meg Comer are married and file a joint tax return. They do not have any children. Matt works as a history professor at a local university and earns a salary of $63,800. Meg works part-time at the same university. She earns $33,100 a year. The couple does not itemize deductions. Other than salary, the Comers’ only other source of income is from the disposition of various capital assets (mostly stocks).

A: What is the Comers’ tax liability for 2019 if they report the following capital gains and losses for the year?

Short-term capital gains $9,020

Short-term capital losses (2,020)

Long-term capital gains 15,030

Long-term capital losses (6,030)

B: What is the Comers’ tax liability for 2019 if they report the following capital gains and losses for the year?

Short-term capital gains $1,500

Short-term capital losses 0

Long-term capital gains 13,040

Long-term capital losses (10,020)
Answer #1

Part A

Total tax liability

$10557

Salary (63800+33100)

96900

Net short term capital gain (short term capital gain – short term capital loss = 9020-2020 )

7000

Net long term capital gain (long term capital gain – long term capital loss = 15030-6030

9000

AGI

112900

Standard deduction (For MFJ 2019)

(24400)

Taxable income

88500

Qualified LTCG (preferentially taxed income)

(9000)

Income taxed at ordinary rates

79500

Ordinary Taxes

9207 (9086+(22%*(79500-78950)))

Preferential tax

1350 (9000*15% as taxable income is above the limit of $78750)

Total tax liability

$10557 (ordinary taxes + preferential taxes)

Part B

Total tax liability

$8492

Salary (63800+33100)

96900

Net short term capital gain (short term capital gain – short term capital loss = 1500-0 )

1500

Net long term capital gain (long term capital gain – long term capital loss = 13040-10020

3020

AGI

101420

Standard deduction (For MFJ 2019)

(24400)

Taxable income

77020

Qualified LTCG (preferentially taxed income)

(3020)

Income taxed at ordinary rates

74000

Ordinary Taxes

8492 (1940+(12%*(74000-19400)))

Preferential tax

0 (taxable income is in the limit of $78750)

Total tax liability

$8492 (ordinary taxes + preferential taxes)